# Burnside type results for fusion categories

### v10

@@ 2. Abelian rational normalizable hypergroups
Let ${(H,\mathcal{B})}$ be a normalized abelian hypergroup and ${\widehat{(H,\mathcal{B})}}$ its dual as defined in the introduction. There is a formula for the dual fusion coefficients ${\widehat{p}}_{k}(i,j)$ given in [Bur21, Proposition 2.1]. For a normalized abelian hypergroup ${(H,\mathcal{B})}$, this is given by:
(2.1) $\widehat{p}_{k}({j_{1}},{j_{2}})=\frac{1}{n_{k}}\bigg{(}\sum_{a\in{\mathcal{B}}}h_{a}\mu_{j_{1}}(a)\mu_{j_{2}}(a)\mu_{k}(a^{*})\bigg{)}$
where the scalars $n_{k}$ are defined below (2.2).
@@
Let ${(H,\mathcal{B})}$ be an abelian normalized hypergroup. We usually let :=${\mathcal{B}}:=\{x_{i}\}_{i\in\mathcal{I}}$, with :=$\mathcal{I}:=\{1,\dots m\}$ and $x_{1}=1$, and say that ${(H,\mathcal{B})}$ has rank $m$. We can simplify the notation $N_{x_{i}x_{j}}^{x_{k}}$ by $N_{i,j}^{k}$. Let $\{x_{i}^{\circ}\}_{i\in\mathcal{I}}$be the linear dual basis of the linear basis ${\mathcal{B}}$ of $H$. Therefore $\langle x_{i}^{\circ},x_{i^{\prime}}\rangle=\delta_{i,i^{\prime}}$, for all $i,i^{\prime}\in\mathcal{I}$. It is easy to see that in this case $\widehat{E}_{i}:=x_{i}^{\circ}\in H^{*}$ is a primitive central idempotent of $H^{*}$ corresponding to the character
@@ 2.1. Function $\tau$ and associative non-degenerate bilinear form
Let $(H,\mathcal{B})$ be any abelian hypergroup. Define $h_{i}:=1/N^{1}_{i,i^{*}}$, for all $i\in\mathcal{I}$. Note that $h_{1}=1$. Then $H$ is a commutative semisimple algebra, and therefore $H\simeq\bigoplus_{j=1}^{m}\mathbb{C}F_{j}$, where $F_{j}$ are the central primitive idempotents of $H$. One can define $\tau\in H^{*}$ with $\tau(x_{i})=\delta_{i,1}$, where as above, $x_{1}=1$. Note that $\tau=x_{1}^{\circ}$ with the above notations. Moreover, in this case
$\begin{array}[]{cccc}(\;,\;)_{\tau}:&H\otimes H&\to&\mathbb{C}\\ &(a,b)&\mapsto&\tau(ab)\end{array}$
is an associative symmetric non-degenerate bilinear form. Thus, one can write
(2.2) $\tau=\sum_{j=1}^{m}\frac{1}{n_{j}}\mu_{j}.$
for some non-zero scalars $n_{j}\in\mathbb{C}^{\times}$. These scalars $n_{j}$ are called the formal codegrees of $(H,\mathcal{B})$, see [Ost15]. Note that $\sum_{j=1}^{m}\frac{1}{n_{j}}=\tau(1)=1$.
From the definition of a hypergroup it follows that $\{h_{i}x_{i}\}_{i\in\mathcal{I}}$ and $\{x_{i^{*}}\}_{i\in\mathcal{I}}$ form a pair of dual bases for $(\;,\;)_{\tau}$. Note that in this case $\{n_{j}F_{j}\}_{i\in\mathcal{I}}$ and $\{F_{j}\}_{i\in\mathcal{I}}$ form also a pair of dual basis for the bilinear form $(\;,\;)_{\tau}$. The existence of the two pairs of dual bases implies that:
(2.3) $\sum_{i=1}^{m}{h_{i}}x_{i}\otimes x_{i}^{*}=\sum_{j=1}^{m}n_{j}F_{j}\otimes F_{j}$
From here it follows that
(2.4) $F_{j}:=\frac{1}{n_{j}}\big{(}\sum_{i=1}^{m}{h_{i}}\mu_{j}(x_{i^{*}})x_{i}\big{)}$
by applying $\operatorname{id}\otimes\mu_{j}$ for the character $\mu_{j}$ corresponding to $F_{j}$. Applying $\mu_{j}\otimes\mu_{k}$ to the above, one obtains the first orthogonality relations
(2.5) $\sum_{i=1}^{m}{h_{i}}\mu_{j}(x_{i})\mu_{k}(x_{i^{*}})=\delta_{j,k}n_{j}$
Since $\mu_{j}(x_{i^{*}})=\overline{\mu_{j}(x_{i})}$, for $j=k$, it follows from here that $n_{j}>0$. By a classical argument, as for finite groups orthogonality, one obtains the second orthogonality relation:
(2.6) $\sum_{j=1}^{m}\frac{1}{n_{j}}\mu_{j}(x_{i})\mu_{j}(x_{l}^{*})=\delta_{i,l}h_{i}^{-1}$
@@ 2.2. Normalizable abelian hypergroups and their duals
Note that hypergroups can be rescaled as follows. Let $y_{i}:=\frac{x_{i}}{\alpha_{{}_{i}}}$, for some non-zero complex numbers $\alpha_{{}_{i}}$ with $\alpha_{1}=1$, $\alpha_{{}_{i^{*}}}=\overline{\alpha_{{}_{i}}}$. Let :=$\mathcal{B}^{\prime}:=\{y_{i}\}_{i\in\mathcal{I}}$. Then it is easy to see that $(H,\mathcal{B}^{\prime})$ is also a hypergroup. However note that $\tau^{\prime}(\frac{x_{i}}{\alpha_{{}_{i}}})=\delta_{i,1}$ is the same functional $\tau$ on $H$ and therefore the formal codegrees $n_{j}$ do not change after rescaling.
@@
Let ${(H,\mathcal{B})}$ be a hypergroup. Then ${(H,\mathcal{B})}$ is normalizable if and only if there is $\mu_{1}\in{\widehat{\mathcal{B}}}$ such that $\mu_{1}(x_{i})\neq 0$ for all $1\leq i\leq m$.
@@
It is easy to see that if $(H,\{\frac{x_{i}}{\alpha_{{}_{i}}}\})$ is normalized after a rescaling, then $\mu_{1}:H\rightarrow H,\;x_{i}\mapsto\alpha_{{}_{i}}$ is a morphism of algebras on $H$. Thus $\mu_{1}\in{\widehat{\mathcal{B}}}$. Conversely, one can check directly that $(H,\{\frac{x_{i}}{\mu_{1}(x_{i})}\})$ is a normalized hypergroup for any morphism $\mu_{1}\in{\widehat{\mathcal{B}}}$ such that $\mu_{1}(x_{i})\neq 0$ for all $i\in\mathcal{I}$. ∎
We denote by ${(H,\mathcal{B},\mu_{1})}$ the normalizable hypergroup obtained by normalizing ${(H,\mathcal{B})}$ by $\mu_{1}\in{\widehat{\mathcal{B}}}$. We also denote by $\overline{{(H,\mathcal{B},\mu_{1})}}$ the normalized hypergroup obtained by rescaling ${(H,\mathcal{B},\mu_{1})}$ as in Lemma 2.2.
@@
Recall from Subsection 1.1 that any abelian normalized hypergroup $H$ admits a dual hypergroup $\widehat{H}$. In the case of $\overline{{(H,\mathcal{B},\mu_{1})}}$, we denote this dual by ${{\widehat{{(H,\mathcal{B},\mu_{1})}}}}$.
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup via $\mu_{1}\in\widehat{\mathcal{B}}$ such that $\mu_{1}(x_{i})\neq 0$ for all $1\leq i\leq m$. From Equation (1.2) it follows that the multiplication on the dual of $\widehat{{(H,\mathcal{B},\mu_{1})}}$ can be written as
(2.7) $[f\star g](\frac{x_{i}}{\mu_{1}(x_{i})})=f(\frac{x_{i}}{\mu_{1}(x_{i})})g(\frac{x_{i}}{\mu_{1}(x_{i})})$
for all $f,g\in H^{*}$. Then it is easy to see that the algebra unit of ${\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is $\mu_{1}$. Moreover, the involution on $^{\widehat{\mathcal{B}}}$, as given in §1.1, becomes:
(2.8) $\mu_{j}^{\#}(x_{i})=\bigg{(}\frac{\mu_{1}(x_{i})}{\overline{\mu_{1}(x_{i}})}\bigg{)}\mu_{j}(x_{i}^{*}).$
@@
$n{(H,\mathcal{B},\mu_{1})}:=\sum_{i=1}^{m}h_{i}|\mu_{1}(x_{i})|^{2}.$
In the abelian case, note that
@@
by Equation (2.5).
We shortly write $n(H)$ when $\mathcal{B}$ and $\mu_{1}$ are implicitly understood. In particular, for a normalized hypergroup one may choose $\mu_{1}(x_{i})=1$ for all $1\leq i\leq m$ and therefore $n(H)=\sum_{i=1}^{m}h_{i}$. For the rest of this paper we use the notation $d_{i}:=\mu_{1}(x_{i})$ for any normalizable hypergroup ${(H,\mathcal{B},\mu_{1})}$.
We say that a normalizable hypergroup ${(H,\mathcal{B},\mu_{1})}$ is weakly-rational if its order is a rational number, i.e. $n{(H,\mathcal{B},\mu_{1})}\in\mathbb{Q}$. Also, we say that ${(H,\mathcal{B},\mu_{1})}$ is weakly-integral if its order is an integer, i.e. $n{(H,\mathcal{B},\mu_{1})}\in\mathbb{Z}$. Note that a fusion ring is weakly-rational if and only if it is weakly-integral since $\mathrm{FPdim}(R)$ is an algebraic number in this case.
@@
Note that the formula from Equation (2.1) can be written for abelian normalizable hypergroups as:
(2.10) $\widehat{p}_{k}({j_{1}},{j_{2}})=\frac{1}{n_{k}}\bigg{(}\sum_{x_{i}\in{\mathcal{B}}}\frac{h_{i}}{d_{i}}\mu_{j_{1}}(x_{i})\mu_{j_{2}}(x_{i})\mu_{k}(x_{i^{*}})\bigg{)}$
@@
By applying (2.10) to ${j_{1}}={j_{2}}^{\#}=j$ and $k=1$, and then (2.5), one obtains (see also [Bur21, Equation (2.10)]),
(2.11) $\widehat{p}_{1}(j,j^{\#})=\frac{n_{j}}{n{(H,\mathcal{B},\mu_{1})}}$
Let ${\widehat{h}}_{j}:=\frac{1}{\widehat{p}_{1}(j,j^{*})}$ be the order of $\mu_{j}$ in ${\widehat{{(H,\mathcal{B},\mu_{1})}}}$. It follows that $n{\widehat{{(H,\mathcal{B},\mu_{1})}}}=\sum_{j=1}^{m}{\widehat{h}}_{j}=n{(H,\mathcal{B},\mu_{1})}(\sum_{j=1}^{m}\frac{1}{n_{j}})=n{(H,\mathcal{B},\mu_{1})}$. ∎
Next proposition holds for any normalized hypergroup, not necessarily abelian.
@@
Suppose that $F_{1}=\sum_{i=1}^{m}\alpha_{{}_{i}}x_{i}$ for some $\alpha_{{}_{i}}\in\mathbb{C}$. Recall that $x_{l}F_{i}=\mu_{i}(x_{l})F_{i}$. Then, $x_{l}F_{1}=F_{1}=\sum_{i=1}^{m}\alpha_{{}_{i}}x_{l}x_{i}$. It follows that $\tau(F_{1})=\alpha_{1}=\tau(x_{l}F_{1})=\alpha_{{}_{l^{*}}}h_{l}^{-1}=\tau(F_{1}x_{l})=\alpha_{l}h_{l^{*}}^{-1}.$ Then $F_{1}=\sum_{i=1}^{m}\alpha_{i}x_{i}={\alpha_{1}}(\sum_{i=1}^{m}{h_{i^{*}}}x_{i})$ and $1=\mu_{1}(F_{1})={\alpha_{1}}(\sum_{i=1}^{m}{h_{i}})$. This implies that $\alpha_{1}=\frac{1}{n(H)}$ and therefore $\alpha_{i}=\frac{h_{i^{*}}}{n(H)}$ for all $i\in\mathcal{I}$. ∎
@@
(2.13) $F_{1}=\frac{1}{n{(H,\mathcal{B},\mu_{1})}}\big{(}\sum_{i=1}^{m}h_{i^{*}}d_{i^{*}}x_{i}\big{)}$
@@
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalized hypergroup. Then ${\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is also an abelian normalized hypergroup and the formal codegrees of ${\widehat{{(H,\mathcal{B},\mu_{1})}}}$ are given by
(2.14) $\widehat{n}_{i}=\frac{n{(H,\mathcal{B},\mu_{1})}}{h_{i}}.$
@@
As already mentioned, [Bur21, Proposition 2.3] shows that ${\widehat{(H,\mathcal{B})}}$ is also an abelian normalized hypergroup.
By Remark 2.1, $x_{i}:H^{*}\rightarrow\mathbb{C}$ are the characters of ${\widehat{(H,\mathcal{B})}}$. By definition of $\widehat{\tau}\in(H^{*})^{*}=H$, $\widehat{\tau}(\mu_{j})=\delta_{j,1}$ where $\mu_{1}$ is the unit of $H^{*}$. Thus $\widehat{\tau}=F_{1}$. Equation (2.2) for $H^{*}$ gives that $\widehat{\tau}=\sum_{i=1}^{m}\frac{1}{\widehat{n}_{i}}x_{i}$. Then Equation (2.12) finishes the proof since $n{\widehat{(H,\mathcal{B})}}=n{(H,\mathcal{B})}$, by Lemma 2.6. ∎
@@
A morphism of normalizable hypergroups $\psi:(H,{\mathcal{B}},\mu_{1})\rightarrow(H^{\prime},{\mathcal{B}}^{\prime},\mu_{1}^{\prime})$ is a morphism of hypergroups with the property that $\mu^{\prime}_{1}\circ f=\mu_{1}$.

### v11

@@ 2. Abelian rational normalizable hypergroups
Let ${(H,\mathcal{B})}$ be a normalized abelian hypergroup and $^{\widehat{(H,\mathcal{B})}}$ its dual as defined in the introduction. There is a formula for the dual fusion coefficients ${\widehat{p}}_{k}(i,j)$ given in [Bur21, Proposition 2.1] by:
(2.1) $\widehat{p}_{k}(i,j)=\frac{1}{n_{k}}\bigg{(}\sum_{a\in{\mathcal{B}}}h_{a}\mu_{i}(a)\mu_{j}(a)\mu_{k}(a^{*})\bigg{)}$
where the scalars $n_{k}$ are defined below (2.2), and $h_{a}:=1/N^{1}_{a,a^{*}}$ (then $h_{1}=1$).
@@
Let ${(H,\mathcal{B})}$ be an abelian normalized hypergroup. We usually let :=${\mathcal{B}}:=\{x_{i}\}_{i\in\mathcal{I}}$, with :=$\mathcal{I}:=\{1,\dots m\}$ and $x_{1}=1$, and say that ${(H,\mathcal{B})}$ has rank $m$. We can simplify the notation $N_{x_{i}x_{j}}^{x_{k}}$ by $N_{i,j}^{k}$, and $h_{x_{i}}$ by $h_{i}$.
Remark 2.2.
Let $\{x_{i}^{\circ}\}_{i\in\mathcal{I}}$ be the linear dual basis of the linear basis ${\mathcal{B}}$ of $H$. Therefore $x_{i}^{\circ}(x_{j})=\delta_{i,j}$, for all $i,j\in\mathcal{I}$. It is easy to see that in this case $\widehat{F}_{i}:=x_{i}^{\circ}\in H^{*}$ is a primitive central idempotent of $H^{*}$ corresponding to the character
@@ 2.1. Function $\tau$ and associative non-degenerate bilinear form
Let $(H,\mathcal{B})$ be any abelian hypergroup. The finite dimensional *-algebra $H$ (so semismple) is commutative, therefore $H\simeq\bigoplus_{i=1}^{m}\mathbb{C}F_{i}$, where $(F_{i})$ are the central primitive idempotents of $H$. We can define $\tau\in H^{*}$ with $\tau(x_{i})=\delta_{i,1}$, where as above, $x_{1}=1$. Note that $\tau=x_{1}^{\circ}$ with the above notations. Moreover, in this case
$\begin{array}[]{cccc}(\;,\;)_{\tau}:&H\times H&\to&\mathbb{C}\\ &(a,b)&\mapsto&\tau(ab)\end{array}$
is an associative symmetric non-degenerate bilinear form. Thus, $\tau(F_{i})$ must be nonzero for all $i$. The nonzero scalars $n_{i}:=\tau(F_{i})^{-1}$ are called the formal codegrees of $(H,\mathcal{B})$, see [Ost15]. We can write
(2.2) $\tau=\sum_{i=1}^{m}\tau(F_{i})\mu_{i}=\sum_{i=1}^{m}\frac{1}{n_{i}}\mu_{i}.$
Note that $\sum_{i\in\mathcal{I}}\frac{1}{n_{i}}=\tau(1)=1$.
From the definition of a hypergroup it follows that $\{h_{i}x_{i}\}_{i\in\mathcal{I}}$ and $\{x_{i^{*}}\}_{i\in\mathcal{I}}$ form a pair of dual bases for $(\;,\;)_{\tau}$. Note that in this case $\{n_{j}F_{j}\}_{j\in\mathcal{I}}$ and $\{F_{j}\}_{j\in\mathcal{I}}$ form also a pair of dual basis for the bilinear form $(\;,\;)_{\tau}$. The existence of the two pairs of dual bases implies that:
(2.3) $\sum_{i\in\mathcal{I}}{h_{i}}x_{i}\otimes x_{i}^{*}=\sum_{j\in\mathcal{I}}n_{j}F_{j}\otimes F_{j}$
By applying $\operatorname{id}\otimes\mu_{j}$, for the character $\mu_{j}$ corresponding to $F_{j}$, it follows that
(2.4) $F_{j}=\frac{1}{n_{j}}\sum_{i\in\mathcal{I}}{h_{i}}\mu_{j}(x_{i^{*}})x_{i}$
Applying $\mu_{j}\otimes\mu_{k}$ to the above, we obtain the first orthogonality relation
(2.5) $\sum_{i\in\mathcal{I}}{h_{i}}\mu_{j}(x_{i})\mu_{k}(x_{i^{*}})=\delta_{j,k}n_{j}.$
Since $\mu_{j}(x_{i^{*}})=\overline{\mu_{j}(x_{i})}$, for $j=k$, it follows from here that $n_{j}>0$. By a classical argument (see, for example, the proof of [LPR23, Theorem 2.4]), as for finite groups orthogonality, we obtain the second orthogonality relation:
(2.6) $\sum_{j\in\mathcal{I}}\frac{1}{n_{j}}\mu_{j}(x_{i})\mu_{j}(x_{l^{*}})=\delta_{i,l}h_{i}^{-1}.$
@@ 2.2. Normalizable abelian hypergroups and their duals
Note that hypergroups can be rescaled as follows. Let $y_{i}:=\frac{x_{i}}{\alpha_{{}_{i}}}$, for some non-zero complex numbers $\alpha_{{}_{i}}$ with $\alpha_{1}=1$, $\alpha_{{}_{i^{*}}}=\overline{\alpha_{{}_{i}}}$. Let :=$\mathcal{B}^{\prime}:=\{y_{i}\}_{i\in\mathcal{I}}$. Then, $(H,\mathcal{B}^{\prime})$ is also a hypergroup, with coefficients rescaled as $\frac{\alpha_{k}}{\alpha_{i}\alpha_{j}}N_{i,j}^{k}$; and then $h_{i}:=(N_{i,i^{*}}^{1})^{-1}$ rescaled as $|\alpha_{i}|^{2}h_{i}$. Consider $\tau^{\prime}$ defined by $\tau^{\prime}(y_{i}):=\delta_{i,1}$, then $\tau^{\prime}(x_{i})=\alpha_{i}\delta_{i,1}=\delta_{i,1}$, because $\alpha_{1}=1$. Therefore $\tau^{\prime}=\tau$ and the formal codegrees $(n_{j})$ are invariant by rescaling.
@@
An hypergroup ${(H,\mathcal{B})}$ is normalizable if and only if there is $\mu_{1}\in{\widehat{\mathcal{B}}}$ such that $\mu_{1}(x_{i})$ nonzero, for all $i\in\mathcal{I}$.
@@
The rescaling $(H,\{\frac{x_{i}}{\alpha_{{}_{i}}}\})$ is normalized if and only if $\sum_{i\in\mathcal{I}}\frac{\alpha_{k}}{\alpha_{i}\alpha_{j}}N_{i,j}^{k}=1$, if and only if $\alpha_{i}\alpha_{j}=\sum_{i\in\mathcal{I}}N_{i,j}^{k}\alpha_{k}$, if and only if $\mu_{1}:x_{i}\mapsto\alpha_{i}$ is an algebra morphism (i.e. an element of $^{\widehat{\mathcal{B}}}$) with $\mu_{1}(x_{i})$ nonzero, for all $i\in\mathcal{I}$. ∎
The normalizable hypergroup ${(H,\mathcal{B})}$ together with $\mu_{1}$ as in Lemma 2.3 is denoted ${(H,\mathcal{B},\mu_{1})}$. The normalized hypergroup $(H,\{\frac{x_{i}}{\mu_{1}(x_{i})}\})$ is denoted $\overline{{(H,\mathcal{B},\mu_{1})}}$.
@@
Recall from §1.1 that any abelian normalized hypergroup $H$ admits a dual hypergroup $H^{*}$. In the case of $\overline{{(H,\mathcal{B},\mu_{1})}}$, we denote this dual by $^{{\widehat{{(H,\mathcal{B},\mu_{1})}}}}$.
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup. From Equation (1.2), it follows that the multiplication on the dual $\widehat{{(H,\mathcal{B},\mu_{1})}}$ can be written as
(2.7) $[f\star g]\left(\frac{x_{i}}{\mu_{1}(x_{i})}\right)=f\left(\frac{x_{i}}{\mu_{1}(x_{i})}\right)g\left(\frac{x_{i}}{\mu_{1}(x_{i})}\right)$
for all $f,g\in H^{*}$. Then it is easy to see that the algebra unit of $^{\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is $\mu_{1}$. Moreover, the involution on $^{\widehat{\mathcal{B}}}$, as given in §1.1, becomes:
(2.8) $\mu_{j^{\#}}(\frac{x_{i}}{\mu_{1}(x_{i})})=\mu_{j}(\frac{x_{i^{*}}}{\overline{\mu_{1}(x_{i}})}).$
@@
$n{(H,\mathcal{B},\mu_{1})}:=\sum_{i\in\mathcal{I}}h_{i}|\mu_{1}(x_{i})|^{2}.$
In the abelian case, the following holds by Equation (2.5):
@@
Proposition 2.6.
Let ${(H,\mathcal{B},\mu_{1})}$ be a normalizable hypergroup. Then the order $n{(H,\mathcal{B},\mu_{1})}$ is invariant under rescaling.
Proof.
Let $x^{\prime}_{i}:=\frac{x_{i}}{\alpha_{{}_{i}}}$, for some non-zero complex numbers $\alpha_{{}_{i}}$ with $\alpha_{1}=1$, $\alpha_{{}_{i^{*}}}=\overline{\alpha_{{}_{i}}}$. Consider $(H,\mathcal{B}^{\prime},\mu_{1})$ the rescaled hypergroup with :=$\mathcal{B}^{\prime}:=\{x^{\prime}_{i}\}_{i\in\mathcal{I}}$. Recall that the order of $x^{\prime}_{i}$ is $h^{\prime}_{i}=|\alpha_{{}_{i}}|^{2}h_{i}>0.$ Therefore
$n(H,\mathcal{B}^{\prime},\mu_{1})=\sum_{i\in\mathcal{I}}h^{\prime}_{i}\mid\mu_{1}(x^{\prime}_{i})\mid^{2}=\sum_{i\in\mathcal{I}}h_{i}\mid\mu_{1}(x_{i})\mid^{2}=n{(H,\mathcal{B},\mu_{1})}.\qed$
We shortly write $n(H)$ when $\mathcal{B}$ and $\mu_{1}$ are implicitly understood. In particular, for a normalized hypergroup we may choose $\mu_{1}(x_{i})=1$ for all $1\leq i\leq m$ and therefore $n(H)=\sum_{i\in\mathcal{I}}h_{i}$. For the rest of this paper we use the notation $d_{i}:=\mu_{1}(x_{i})$ for any normalizable hypergroup ${(H,\mathcal{B},\mu_{1})}$.
We say that a normalizable hypergroup ${(H,\mathcal{B},\mu_{1})}$ is weakly-rational if its order is a rational number, i.e. $n{(H,\mathcal{B},\mu_{1})}\in\mathbb{Q}$. Also, we say that ${(H,\mathcal{B},\mu_{1})}$ is weakly-integral if its order is an integer, i.e. $n{(H,\mathcal{B},\mu_{1})}\in\mathbb{Z}$. Note that a fusion ring is weakly-rational if and only if it is weakly-integral since $\mathrm{FPdim}(R)$ is an algebraic integer in this case.
@@
Note that the formula from Equation (2.1) can be written for abelian normalizable hypergroups as follows (using the fact that $h_{l}$ rescales as $|d_{l}|^{2}h_{l}$):
(2.10) $\widehat{p}_{k}(i,j)=\frac{1}{n_{k}}\bigg{(}\sum_{x_{l}\in{\mathcal{B}}}\frac{h_{l}}{d_{l}}\mu_{i}(x_{l})\mu_{j}(x_{l})\mu_{k}(x_{l^{*}})\bigg{)}$
@@
By applying (2.10) to $i=j^{\#}$ and $k=1$, and then (2.5), we obtain (see also [Bur21, Equation (2.10)]),
(2.11) $\widehat{p}_{1}(i,i^{\#})=\frac{n_{i}}{n{(H,\mathcal{B},\mu_{1})}}$
Let ${\widehat{h}}_{i}:=\frac{1}{\widehat{p}_{1}(i,i^{\#})}$ be the order of $\mu_{i}$ in $^{\widehat{{(H,\mathcal{B},\mu_{1})}}}$. It follows that $n{\widehat{{(H,\mathcal{B},\mu_{1})}}}=\sum_{i\in\mathcal{I}}{\widehat{h}}_{i}=n{(H,\mathcal{B},\mu_{1})}(\sum_{i\in\mathcal{I}}\frac{1}{n_{i}})=n{(H,\mathcal{B},\mu_{1})}$. ∎
The next proposition holds for any normalized hypergroup, not necessarily abelian.
@@
Suppose that $F_{1}=\sum_{i\in\mathcal{I}}\alpha_{{}_{i}}x_{i}$ for some $\alpha_{{}_{i}}\in\mathbb{C}$. Recall that $x_{l}F_{i}=\mu_{i}(x_{l})F_{i}$. Then, $x_{l}F_{1}=F_{1}=\sum_{i\in\mathcal{I}}\alpha_{{}_{i}}x_{l}x_{i}$. It follows that $\tau(F_{1})=\alpha_{1}=\tau(x_{l}F_{1})=\alpha_{l^{*}}h_{l}^{-1}.$ Thus $\alpha_{l^{*}}=h_{l}\alpha_{1}$, so $\alpha_{l}=h_{l^{*}}\alpha_{1}$, $F_{1}=\sum_{i\in\mathcal{I}}\alpha_{i}x_{i}={\alpha_{1}}(\sum_{i\in\mathcal{I}}{h_{i^{*}}}x_{i})$ and $1=\mu_{1}(F_{1})={\alpha_{1}}(\sum_{i\in\mathcal{I}}{h_{i}})$. This implies that $\alpha_{1}=\frac{1}{n(H)}$ and therefore $\alpha_{i}=\frac{h_{i^{*}}}{n(H)}$ for all $i\in\mathcal{I}$. ∎
@@
(2.13) $F_{1}=\frac{1}{n{(H,\mathcal{B},\mu_{1})}}\big{(}\sum_{i\in\mathcal{I}}h_{i^{*}}d_{i^{*}}x_{i}\big{)}$
Remark 2.11.
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian (rational) normalizable hypergroup. In order to find the central primitive idempotents of $H^{*}$, we can apply Remark 2.2 after normalization. If $x_{i}^{\circ}\in H^{*}$ are defined as above by $\langle x_{i}^{\circ},x_{i^{\prime}}\rangle=\delta_{i,i^{\prime}}$, then $\widehat{F}_{i}:=d_{i}x_{i}^{\circ}\in H^{*}$ are the primitive central idempotents of $H^{*}$ corresponding to the character
$\begin{array}[]{cccc}\widetilde{\omega}_{i}=\operatorname{ev}_{\overline{x}_{i}}:&H^{*}&\to&\mathbb{C}\\ &\mu_{j}&\mapsto&\mu_{j}(\overline{x}_{i}),\end{array}$
where $\overline{x}_{i}=\frac{x_{i}}{d_{i}}$. By identifying $H^{**}$ with $H$, we get that $\operatorname{ev}_{\overline{x}_{i}}=\overline{x}_{i}$.
Recall that $\overline{x}_{1}=x_{1}$.
@@
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup. Then $^{\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is an abelian normalized hypergroup whose formal codegrees are
(2.14) $\widehat{n}_{i}=\frac{n{(H,\mathcal{B},\mu_{1})}}{h_{i}|d_{i}|^{2}}.$
@@
As already mentioned, [Bur21, Proposition 2.3] shows that $^{\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is an abelian normalized hypergroup.
Let $\overline{x}_{i}=\frac{x_{i}}{d_{i}}$ be the normalization of $x_{i}$. By Remark 2.11, $\overline{x}_{i}:H^{*}\rightarrow\mathbb{C}$ are the characters of $^{\widehat{{(H,\mathcal{B},\mu_{1})}}}$. By definition of $\widehat{\tau}\in(H^{*})^{*}=H$, $\widehat{\tau}(\mu_{j})=\delta_{j,1}$ as $\mu_{1}$ is the unit of $H^{*}$. Thus $\widehat{\tau}=F_{1}$. Equation (2.2) for $H^{*}$ gives that $\widehat{\tau}=\sum_{i\in\mathcal{I}}\frac{1}{\widehat{n}_{i}}\overline{x}_{i}$. So, by Equation (2.13),
$\frac{1}{n{(H,\mathcal{B},\mu_{1})}}\big{(}\sum_{i\in\mathcal{I}}h_{i^{*}}d_{i^{*}}x_{i}\big{)}=F_{1}=\widehat{\tau}=\sum_{i\in\mathcal{I}}\frac{1}{\widehat{n}_{i}}\frac{x_{i}}{d_{i}}.$
The result follows by $h_{i^{*}}=(N_{i^{*},i}^{1})^{-1}=(N_{i,i^{*}}^{1})^{-1}=h_{i}$, in the abelian case. ∎
@@
A morphism of normalizable hypergroups $\psi:(H,{\mathcal{B}},\mu_{1})\rightarrow(H^{\prime},{\mathcal{B}}^{\prime},\mu_{1}^{\prime})$ is a morphism of hypergroups with the property that $\mu^{\prime}_{1}\circ\psi=\mu_{1}$.